# The evolution of cooperation with Q-learning: the impact of information perception

### v1

@@ IV.3 Stage III — Cooperation reestablishment
In summary, we investigate the evolution of the iterated prisoner’s dilemma game under three distinct information scenarios within the framework of reinforcement learning. Diverging from existing research on cooperation within this paradigm, our primary focus is on the impact of different information on the evolution of cooperation. Specifically, we examined three simplest information settings. Our findings reveal that the information structure significantly influences the evolution of cooperation. In the two symmetric scenarios, the direct association between the actions and states is more conducive to the emergence of cooperation. Remarkably, cooperation is more likely to emerge in the asymmetric scenario and does so more rapidly. The evolutionary dynamics exhibit characteristics of a first-order-like phase transition, with the average cooperation preference transitions between mutual cooperation and mutual defection. Further mechanism analysis elucidates how cooperation emerges, breaks down, and reconstructs. The basin of attraction for the two stable states are identified for some typical dilemma intensities.
The majority of existing research on cooperation primarily focuses on how cooperation emerges and is sustained Roca et al. (2009b); Sigmund (2010). There are very few work discussing how cooperation breaks down and the conditions necessary for cooperative reconstruction Ding et al. (2023). Our study provides insights into the establishment of long-term cooperation, suggesting that a moderate and tolerant understanding may prolong cooperative relationships. However, individual tolerance towards others is often limited, and relentless exploitation can lead to the collapse of cooperative ties. These observations are in line with our daily experience. Additionally, the reconstruction of cooperation is not straightforward, during this process, the player previously engaged in exploitation often finds itself in a disadvantaged position.
Certainly, our work serves as an initial effort to understand the impact of information perception on cooperation within the RL framework. We only consider three simplest information structures within two-body scenarios. In the real world, however, due to the diversity of personal and society factors Henrich et al. (2004), people may grab very different parts of information. Furthermore, people are connected within complicated social relationships which are often depicted as complex networks Newman (2018), and their information perception may also influence with each other through the networks. These complexities are far more complex beyond what we studied here and require further systematic studies in the future.
VII Acknowledgments
@@
Appendix A Appendix A: Asymmetric information causes imbalanced Q-value evolution between two players
A.1 Within sub-stage III of stage-I
@@
There the system in a (D,C)$\leftrightarrow$(C,D) cyclic state. Individual M cooperates in state D and defects in state C, while individual Y cooperates in state C and defects in state D.
@@
For individual M: exploratory cooperation in state C shifts the action preference from (D,C)$\leftrightarrow$(C,D) to (C,C)$\leftrightarrow$(D,C), resulting in an immediate mutual cooperation payoff of $R=1$ and an increase in $Q_{C,C}^{M}$. This exploratory behavior also drives the system to state C, yielding a temptation value of $T=1.2$ under the (D,C) preference, thus accelerating the increase in $Q_{C,C}^{M}$.
@@
For individual Y: exploratory cooperation in state D shifts the action preference from (D,C)$\leftrightarrow$(C,D) to (C,C)$\leftrightarrow$(D,C), resulting in an immediate mutual cooperation payoff of $R=1$ and an increase in $Q_{D,C}^{Y}$. However, since individual Y cannot directly alter the system’s state, it continues along its previous trajectory into state C, where under the (D,C) action preference, it receives the payoff for sucker, $S=-0.2$, without the additional incentive seen in individual M.
A.2 Within sub-stage IV of stage-I
There the system in a (C,C) state, both individuals choose to cooperate in state C.
@@ A.2 Within sub-stage IV of stage-I
For individual M: exploratory defection in state C drives the system to state C, shifting the action preference combination from (C,C) to (D,C)$\leftrightarrow$(C,D), then back to (C,C) [consistent with the process in sub-stage III]. Then, an immediate temptation payoff of $T=1.2$ is obtained and $Q_{C,D}^{M}$ is increased.
@@
For individual Y: exploratory defection in state C shifts the action preference combination from (C,C) to (C,D), then back to (C,C). This results in an immediate temptation payoff of $T=1.2$, leading to an increase in $Q_{C,D}^{Y}$. However, unlike individual M, Y cannot alter the system’s state, thus bypassing the process of reverting to sub-stage III, consequently accelerating the increase in $Q_{C,D}^{Y}$.
A.3 Within sub-stage II of stage-III
@@
There the system in a (C,C)$\leftrightarrow$(D,D) cyclic state. Both individuals choose to cooperate in state D and defect in state C.
@@
For individual M: exploratory defection in state D shifts the action preference from (C,C)$\leftrightarrow$(D,D) to (D,C)$\leftrightarrow$(C,C), resulting in an immediate temptation payoff of $T=1.2$ and an increase in $Q_{D,D}^{M}$. This exploratory behavior also drives the system to state D, yielding a mutual cooperation payoff of $R=1$ under the (C,C) preference, thus accelerating the increase in $Q_{D,D}^{M}$.
@@
For individual Y: exploratory defection in state D shifts the action preference from (C,C)$\leftrightarrow$(D,D) to (C,D)$\leftrightarrow$(D,D), resulting in an immediate temptation payoff of $T=1.2$ and an increase in $Q_{D,D}^{Y}$. However, since individual Y cannot alter the system’s state, it continues along its previous trajectory into state C, where under the (D,D) action preference, it receives the payoff for punishment, $P=0$, without the additional incentive seen in individual M.
Appendix B Appendix B: mechanism analysis in Scheme I
@@
In Scheme I, the states of both parties directly depend on their opponent’s action information. However, there is a key issue in maintaining mutual cooperation: both parties must always choose to cooperate in state C. Once one party shifts from cooperation to betrayal, the system will enter a stage of mutual exploitation. Because actions under this scheme cannot directly change the state, it is impossible to implement punishment-like policies similar to those in Scheme II and Scheme III. Therefore, when exploratory betrayal behaviors accumulate advantages over time, the cooperation conditions will no longer be met, ultimately leading the system into a state of mutual betrayal. As can be seen from the figure 11, cooperation fails to emerge and be sustained.
For more details, we denote two individuals as $i={Y_{1},Y_{2}}$, who consider their opponent’s action information. The values of $Q_{s_{l},a_{m}}$ and $\Delta Q_{s_{l}}^{i}$ are labeled as in the text. Even starting from an initial condition of full cooperation, i.e., $Q_{C,C}^{Y_{1},Y_{2}}>Q_{C,D}^{Y_{1},Y_{2}}$ and $Q_{D,C}^{Y_{1},Y_{2}}>Q_{D,D}^{Y_{1},Y_{2}}$, occasional exploratory choices of betrayal by both parties will lead to an increase in $Q_{C,D}^{Y_{1},Y_{2}}$.
After the advantage of betrayal accumulates over time, satisfying $Q_{C,D}^{Y_{1}/Y_{2}}>Q_{C,C}^{Y_{1}/Y_{2}}$, one individual switches from cooperation to betrayal in state C. This initiates a continuous exploitation process (C,D)--(C,D)--(C,D). As continuous exploitation causes the other individual¡¯s tendency to cooperate in state D to decline, they eventually switch to betrayal in state D, leading to another continuous exploitation process with roles reversed (D,C)--(D,C)--(D,C). This process ultimately results in both individuals having no inclination to choose cooperation in either state, leading to the tragedy of total betrayal.
Appendix C Appendix C: mechanism analysis in Scheme II
@@
Figure 13: Dynamical evolution process in Scheme II. The figures show the evolution of action combination preferences, and the temporal evolution of $\Delta Q_{s_{l}}^{M_{1},M_{2}}$. Here, the action combination (C,C)(D,D) indicates that both individuals choose to defect in state C and to cooperate in state D. An exploratory action by one party can disrupt this synchronization, leading to (C,D)(D,C), while still maintaining the same action preferences. Thus, (C,C)(D,D) (C,D)(D,C) represent the alternation between synchronized and unsynchronized states under the influence of exploratory actions. The action combinations (C,D)(D,D) and (D,C)(D,D) indicate that one individual chooses to defect in state D, while the other defects in state C and cooperates in state D. Therefore, (C,D)(D,D) (D,C)(D,D) represent this process occurring sequentially and swapping the positions of the two individuals. This can be viewed as an alternating exploitation and punishment process. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@
    Stage-I: Mutual betrayal.
@@
    Stage-II: Breaking away from mutual betrayal.
@@
    Stage-III: Establishing and maintaining mutual cooperation.
Here, $i=\{M_{1},M_{2}\}$ respectively labels the two players, who consider their own action information, the values of $Q_{s_{l},a_{m}}$ and $\Delta Q_{s_{l}}^{i}$ are labeled in the same way as they are in the text. We initiate the study from initial conditions far from cooperation, i.e., $Q_{CC}^{M_{1},M_{2}}<Q_{CD}^{M_{1},M_{2}}$ and $Q_{DC}^{M_{1},M_{2}}<Q_{DD}^{M_{1},M_{2}}$, and analyze the mechanism in stages:
@@ C.1 Stage I — Mutual betrayal
During this stage, mutual defection (D, D) does not yield any payoffs for either party, leading to an increase in $\Delta Q_{D}^{M_{1},M_{2}}$. Intermittent decreases occur due to exploratory cooperation. When $\Delta Q_{D}^{M_{1}/M_{2}}>0$, his/her preference shifts from defection (D) to cooperation (C). However, unilateral cooperation merely strengthens the other player’s preference for defection, as the only perceivable change is an increased payoff for maintaining the original action. Therefore, breaking the (D, D) preference through a unilateral shift is challenging. As shown in Fig. 13(a), $\Delta Q_{D}^{M_{0},M_{1}}$ fluctuates but remains consistently below 0.
@@ C.2 Stage II — Breaking away from mutual betrayal
Sub-stage ISimultaneous cooperative exploration breaks mutual defection, triggers (C,C) $\leftrightarrow$ (D,D) cyclic state.
When both individuals simultaneously engage in exploratory cooperative behavior, they achieve positive payoffs $R$, which leads to a continuous increase in $Q_{D,C}^{M_{1},M_{2}}$. This results in $\Delta Q_{D}^{M_{1},M_{2}}>0$ and a reversal in preference [Fig. 13(b-2)]. The system then cycles between (C,C) $\leftrightarrow$ (D,D). Subsequent exploratory behavior disrupts this synchronization, forming the combinations (C,D) $\leftrightarrow$ (D,C). Consequently, synchronization and asynchronization alternate [Fig. 13(b-1)], with both parties choose to cooperate in state D and defect in state C. However, this action preference combination fails to persist with weak exploration.
@@
Within the above action preference combination, both individuals’ exploratory defection in state D leads to intermittent increases in $Q_{D,D}^{M_{1},M_{2}}$. When $Q_{D,D}^{M_{1}/M_{2}}\textgreater Q_{D,C}^{M_{1}/M_{2}}$, the system forms the action preference combination (C,D) $\leftrightarrow$ (D,D), with (D,D) occurring more frequently. This can be viewed as a process where one party punishes the other, resulting in an increasing trend in $\Delta Q_{C}^{M_{1}/M_{2}}$[Fig. 13(c)]. When $\Delta Q_{C}^{M_{1}/M_{2}}$ >00, the action preference in state C shifts towards cooperation. As indicated by the rectangular dotted boxes, when this process occurs sequentially for both individuals, their preference for defection in state C transitions to cooperation, establishing a (C,C) positive feedback loop.

### v2

@@ VI Discussion
In summary, we explore the evolution of cooperation in the iterated prisoner’s dilemma game under three distinct information scenarios within the reinforcement learning (RL) framework. Unlike existing studies, we focus on how different information perceptions influence cooperation dynamics. Our findings demonstrate that information structure plays a critical role: in symmetric scenarios, direct action-state associations foster cooperation, while the asymmetric scenario promotes faster and more robust cooperation emergence. The evolutionary dynamics exhibit first-order-like phase transitions, with cooperation preference oscillating between mutual cooperation and defection. Mechanism analysis reveals the processes of cooperation emergence, breakdown, and reconstruction, alongside identifying basins of attraction for stable states at specific dilemma intensities.
While most research focuses on the emergence and maintenance of cooperation Roca, Cuesta, and Sánchez (2009); Sigmund (2010), few address its breakdown and reconstruction Ding et al. (2023). Our study highlights that moderate tolerance can sustain cooperation, but excessive exploitation risks its collapse, aligning with real-world observations. Rebuilding cooperation is challenging, often leaving exploiters at a disadvantage.
This work is an initial step in understanding information perception’s role in cooperation within RL. We limit our analysis to three simple information structures in two-player scenarios, but real-world complexities—such as diverse personal and societal factors Henrich et al. (2004) and intricate social networks Newman (2018)—warrant further investigation. Additionally, integrating moral preference hypotheses Capraro and Perc (2021) with RL to better simulate decision-making presents a promising future direction.
Acknowledgements.
@@
Appendix A Asymmetric information causes imbalanced Q-value evolution between two players
A.1 Within sub-stage iii of stage i
@@
There the system in a $(C,D)_{D}$leftrightarrow $(D,C)_{C}$ cyclic state. Individual M cooperates in state D and defects in state C, while individual Y cooperates in state C and defects in state D.
@@
For individual M: exploratory cooperation in state C shifts the action preference from $(C,D)_{D}$leftrightarrow $(D,C)_{C}$ to $(C,D)_{D}$leftrightarrow $(C,C)_{C}$, resulting in an immediate mutual cooperation payoff of $R=1$ and an increase in $Q_{C,C}^{M}$. This exploratory behavior also drives the system to state C, yielding a temptation value of $T=1.2$ under the $(D,C)_{C}$ preference, thus accelerating the increase in $Q_{C,C}^{M}$.
@@
For individual Y: exploratory cooperation in state D shifts the action preference from $(C,D)_{D}$leftrightarrow $(D,C)_{C}$ to $(C,C)_{D}$leftrightarrow $(D,C)_{C}$, resulting in an immediate mutual cooperation payoff of $R=1$ and an increase in $Q_{D,C}^{Y}$. However, since individual Y cannot directly alter the system’s state, it continues along its previous trajectory into state C, where under the $(D,C)_{C}$ action preference, it receives the payoff for sucker, $S=-0.2$, without the additional incentive seen in individual M.
A.2 Within sub-stage iv of stage i
There the system in a $(C,C)_{C}$ state, both individuals choose to cooperate in state C.
@@ A.2 Within sub-stage iv of stage i
For individual M: exploratory defection in state C drives the system to state D, shifting the action preference combination from $(C,C)_{C}$ to $(D,C)_{C}$leftrightarrow $(C,D)_{D}$, then back to $(C,C)_{C}$ [consistent with the process in sub-stage iii]. Then, an immediate temptation payoff of $T=1.2$ is obtained and $Q_{C,D}^{M}$ is increased.
@@
For individual Y: exploratory defection in state C shifts the action preference combination from $(C,C)_{C}$ to $(C,D)_{C}$, then back to $(C,C)_{C}$. This results in an immediate temptation payoff of $T=1.2$, leading to an increase in $Q_{C,D}^{Y}$. However, unlike individual M, Y cannot alter the system’s state, thus bypassing the process of reverting to sub-stage iii, consequently accelerating the increase in $Q_{C,D}^{Y}$.
A.3 Within sub-stage ii of stage iii
@@
There the system in a $(C,C)_{D}$leftrightarrow $(D,D)_{C}$ cyclic state. Both individuals choose to cooperate in state D and defect in state C.
@@
For individual M: exploratory defection in state D shifts the action preference from $(C,C)_{D}$leftrightarrow $(D,D)_{C}$ to $(D,C)_{D}$leftrightarrow $(C,C)_{D}$, resulting in an immediate temptation payoff of $T=1.2$ and an increase in $Q_{D,D}^{M}$. This exploratory behavior also drives the system to state D, yielding a mutual cooperation payoff of $R=1$ under the $(C,C)_{D}$ preference, thus accelerating the increase in $Q_{D,D}^{M}$.
@@
For individual Y: exploratory defection in state D shifts the action preference from $(C,C)_{D}$leftrightarrow $(D,D)_{C}$ to $(C,D)_{D}$leftrightarrow $(D,D)_{C}$, resulting in an immediate temptation payoff of $T=1.2$ and an increase in $Q_{D,D}^{Y}$. However, since individual Y cannot alter the system’s state, it continues along its previous trajectory into state C, where under the $(D,D)_{C}$ action preference, it receives the payoff for punishment, $P=0$, without the additional incentive seen in individual M.
Appendix B Mechanism analysis in Scheme I
@@
In Scheme I, both individuals focus on the opponent’s actions, resembling the Tit for Tat (TFT) strategy but falling short of fully implementing it. A key limitation is the difficulty in establishing and maintaining a “cooperation-cooperation” pattern. Moreover, since individuals cannot directly determine the state, the system lacks the ability to enforce punishment-like strategies, as seen in Schemes II and III. Even when starting from TFT-like initial conditions, cooperation proves unsustainable. Occasional misunderstandings gradually erode the tendency to cooperate in state C, ultimately leading to mutual defection.
The underlying evolutionary mechanism involves a key issue in maintaining mutual cooperation: both parties must always choose to cooperate in state C. Once one party shifts from cooperation to betrayal, the system will enter a stage of mutual exploitation. When exploratory betrayal behaviors accumulate advantages over time, the cooperation conditions will no longer be met, ultimately leading the system into a state of mutual betrayal. As can be seen from the Fig. 11, cooperation fails to emerge and be sustained.
For more details, we denote two individuals as $i=\{Y_{1},Y_{2}\}$, who consider their opponent’s action information. The values of $Q_{s_{l},a_{m}}$ and $\Delta Q_{s_{l}}^{i}$ are labeled as in the text. Even starting from an initial condition of full cooperation, i.e., $Q_{C,C}^{Y_{1},Y_{2}}>Q_{C,D}^{Y_{1},Y_{2}}$ and $Q_{D,C}^{Y_{1},Y_{2}}>Q_{D,D}^{Y_{1},Y_{2}}$, occasional exploratory choices of betrayal by both parties will lead to an increase in $Q_{C,D}^{Y_{1},Y_{2}}$.
After the advantage of betrayal accumulates over time, satisfying $Q_{C,D}^{Y_{1}/Y_{2}}>Q_{C,C}^{Y_{1}/Y_{2}}$, one individual switches from cooperation to betrayal in state C. This initiates a continuous exploitation process $(C_{D},D_{C})$--$(C_{D},D_{C})$--$(C_{D},D_{C})$. As continuous exploitation causes the other individual’s tendency to cooperate in state D to decline, they eventually switch to betrayal in state D, leading to another continuous exploitation process with roles reversed $(D_{C},C_{D})$--$(D_{C},C_{D})$--$(D_{C},C_{D})$. This process ultimately results in both individuals having no inclination to choose cooperation in either state, leading to the tragedy of total betrayal.
Appendix C Mechanism analysis in Scheme II
Figure 13: Dynamical evolution process in Scheme II. The figures show the evolution of action combination preferences, and the temporal evolution of $\Delta Q_{s_{l}}^{M_{1},M_{2}}$. Here, the action combination $(C,C)_{D}$$(D,D)_{C}$ indicates that both individuals choose to defect in state C and to cooperate in state D. An exploratory action by one party can disrupt this synchronization, leading to $(C_{D}$,$D_{C})$$(D_{C}$,$C_{D})$, while still maintaining the same action preferences. Thus, $(C,C)_{D}$$(D,D)_{C}$ $(C_{D}$,$D_{C})$$(D_{C}$,$C_{D})$ represent the alternation between synchronized and unsynchronized states under the influence of exploratory actions. The action combinations $(C,D)_{D}$ and $(D,C)_{D}$ indicate that in state D, one individual cooperates while the other defects. Therefore, $(C,D)_{D}(D,D)_{D}$ $(D,C)_{D}(D,D)_{D}$ represent this process occurring sequentially and swapping the positions of the two individuals. This can be viewed as an alternating exploitation and punishment process. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@ Appendix C Mechanism analysis in Scheme II
    Stage i: Mutual betrayal.
@@
    Stage ii: Breaking away from mutual betrayal.
@@
    Stage iii: Establishing and maintaining mutual cooperation.
Here, $i=\{M_{1},M_{2}\}$ respectively labels the two players, who consider their own action information, the values of $Q_{s_{l},a_{m}}$ and $\Delta Q_{s_{l}}^{i}$ are labeled in the same way as they are in the text. We initiate the study from initial conditions far from cooperation, i.e., $Q_{C,C}^{M_{1},M_{2}}<Q_{C,D}^{M_{1},M_{2}}$ and $Q_{D,C}^{M_{1},M_{2}}<Q_{D,D}^{M_{1},M_{2}}$, and analyze the mechanism in stages.
@@ C.1 Stage i — Mutual betrayal
During this stage, mutual defection $(D,D)_{D}$ does not yield any payoffs for either party, leading to an increase in $\Delta Q_{D}^{M_{1},M_{2}}$. Intermittent decreases occur due to exploratory cooperation. When $\Delta Q_{D}^{M_{1}/M_{2}}>0$, his/her preference shifts from defection (D) to cooperation (C). However, unilateral cooperation merely strengthens the other player’s preference for defection, as the only perceivable change is an increased payoff for maintaining the original action. Therefore, breaking the $(D,D)_{D}$ preference through a unilateral shift is challenging. As shown in Fig. 13(a), $\Delta Q_{D}^{M_{0},M_{1}}$ fluctuates but remains consistently below 0.
@@ C.2 Stage ii — Breaking away from mutual betrayal
Sub-stage iSimultaneous cooperative exploration breaks mutual defection, triggers $(C,C)_{D}$leftrightarrow $(D,D)_{C}$ cyclic state.
When both individuals simultaneously engage in exploratory cooperative behavior, they achieve positive payoffs $R$, which leads to a continuous increase in $Q_{D,C}^{M_{1},M_{2}}$. This results in $\Delta Q_{D}^{M_{1},M_{2}}>0$ and a reversal in preference [Fig. 13(b-2)]. The system then cycles between $(C,C)_{D}$leftrightarrow $(D,D)_{C}$. Subsequent exploratory behavior disrupts this synchronization, forming the combinations $(C_{D},D_{C})$leftrightarrow $(D_{C},C_{D})$. Consequently, synchronization and asynchronization alternate [Fig. 13(b-1)], with both parties choose to cooperate in state D and defect in state C. However, this action preference combination fails to persist with weak exploration.
@@
Within the above action preference combination, both individuals’ exploratory defection in state D leads to intermittent increases in $Q_{D,D}^{M_{1},M_{2}}$. When $Q_{D,D}^{M_{1}/M_{2}}\textgreater Q_{D,C}^{M_{1}/M_{2}}$, the system forms the action preference combination $(C,D)_{D}$leftrightarrow $(D_{C},D_{D})$, with $(D,D)_{D}$ occurring more frequently. This can be viewed as a process where one party punishes the other, resulting in an increasing trend in $\Delta Q_{C}^{M_{1}/M_{2}}$[Fig. 13(c)]. When $\Delta Q_{C}^{M_{1}/M_{2}}$ >00, the action preference in state C shifts towards cooperation. As indicated by the rectangular dotted boxes, when this process occurs sequentially for both individuals, their preference for defection in state C transitions to cooperation, establishing a $(C,C)_{C}$ positive feedback loop.